OpenStudy (anonymous):

Given the functions f(n) = 11 and g(n) = (three over four)n − 1, combine them to create a geometric sequence, an, and solve for the 9th term.

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

\[f(n) = 11 ~~~g(n) = \frac {3}{4} n-1\] I'm not too sure how you would combine them to make a sequence..

OpenStudy (shamil98):

You would just add them right?

OpenStudy (anonymous):

an = 11(three fourths)n − 1; a9 ≈ 1.101

OpenStudy (anonymous):

this seems like the most accurate option

OpenStudy (shamil98):

let h(n) = f(n) + g(n) \[h(n) = \frac{ 3n-3 }{ 4 }+11\]

OpenStudy (shamil98):

\[h(9) = \frac{ 3(9-1) }{ 4 } + 11\] \[h(9) = \frac{ 24 }{ 4 } + 11 = 6 + 11 = 17\]

OpenStudy (shamil98):

This is just my guess.

OpenStudy (anonymous):

Is h(n) a geometric sequence? @shamil98 and Why?

OpenStudy (anonymous):

an = (11 • three fourths)n − 1; a9 ≈ 24.301 an = 11(three fourths)n − 1; a9 ≈ 1.101 an = 11 + (three fourths)n − 1; a9 ≈ 11.100 an = 11 − (three fourths)n − 1; a9 ≈ 9.900 these are the options

OpenStudy (shamil98):

Oh, h(n) would be an arithmetic sequence, the difference between terms is .75.

OpenStudy (anonymous):

They are asking to get a geometric sequence

OpenStudy (anonymous):

@eliassaab Which one looks like the right one to you?

OpenStudy (shamil98):

Yeah, so you can cross out options C and D then..

OpenStudy (anonymous):

i think its B

OpenStudy (shamil98):

I think it would be option B.

OpenStudy (anonymous):

What are the options C and D?

OpenStudy (anonymous):

an = (11 • three fourths)n − 1; a9 ≈ 24.301 an = 11(three fourths)n − 1; a9 ≈ 1.101 an = 11 + (three fourths)n − 1; a9 ≈ 11.100 an = 11 − (three fourths)n − 1; a9 ≈ 9.900

OpenStudy (shamil98):

Is it (3/4n) -1 or (3/4) * n-1 ?

OpenStudy (anonymous):

thanks @shamil98 B was correct!!

OpenStudy (anonymous):

B is not a geometric series.

OpenStudy (shamil98):

The question is asking for a sequence not a series.

OpenStudy (anonymous):

'Twas corrrect. :)